Circular No. 8689

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

(42355) 2002 CR_{46}

K. S. Noll, Space Telescope Science Institute; W. M. Grundy, Lowell Observatory; D. C. Stephens, Johns Hopkins University; and H. F. Levison, Southwest Research Institute, report the detection of a binary companion to this centaur-type object (MPECs 2002-C84, 2002-H53; MPS 50182) when at 16.68 AU from the earth on Jan. 20.410–20.426 UT with the High Resolution Camera (+ clear filter) of the Advanced Camera for Surveys (HRC/ACS) on the Hubble Space Telescope as part of an ongoing observing program. The minor planet was observed with one 300-sec exposure at each of four dithered positions on the detector. Two components are clearly resolved in each image and in the co-added image. The fainter component is ≈ 1.2 magnitudes fainter than the brighter of the two. The two components are separated in the images by an angular distance of 0′′.11 ± 0′′.01, with the fainter component at p.a. 226^{o} as measured from the primary. The spacecraft tracked the motion of (42355) as it moved at an average rate of 0′′.085/min; the relative position of the two components remained the same during the orbit. The projected separation of the objects in the plane of the sky is 1330 ± 130 km.

(60458) 2000 CM_{114}

K. S. Noll, W. M. Grundy, H. F. Levison, and D. C. Stephens also report the detection of a binary companion to this transneptunian minor planet (MPEC 2000-J45, 2002-D20; MPO 44436), obtained as above on Jan. 16.276–16.292 UT, when it was more than 41 AU from the earth. Two components are clearly resolved in each image and in the co-added image. The two components are separated in the images by an angular distance of 0′′.073 ± 0′′.006 with the fainter component at a position angle of 155^{o}.2 as measured from the primary. The projected separation of the objects in the plane of the sky is 2200 ± 200 km. The fainter component is ≈ 0.5 magnitude fainter than the brighter of the two. The spacecraft tracked the motion of (60458) as it moved at an average rate of 0′′.047/min; the relative position of the two components remained the same throughout the orbit.

SUPERNOVAE 2005my, 2006O, 2006C, 2006J, 2006Y, 2006ai, 2006ao

Supernovae 2005my (cf. IAUC 8655), 2006O (IAUC 8662), 2006C (IAUC 8657), 2006J (IAUC 8661), 2006Y (IAUC 8668), 2006ai (IAUC 8674), and 2006ao (IAUC 8681) all appear to be type-II supernovae; details appear on CBETs 358, 378, 380, 417, and 422.

2006 March 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green